OpenStudy (anonymous):

Solve the percent problem for the unknown value of p. What percent of 120 is 30?

OpenStudy (ash2326):

let it be x % so120*x%=30 120*x/100=30 120x=3000 x=3000/120=25 so it's 25%

OpenStudy (anonymous):

im lerning more here than i did in class today

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

learning